### v1

# Privacy-Aware Semantic Cache for Large Language Models

Abstract.
Large Language Models (LLMs) like ChatGPT, Google Bard, Claude, and Llama 2 have revolutionized natural language processing and search engine dynamics. However, these models incur exceptionally high computational costs. For instance, GPT-3 consists of 175 billion parameters and inference on these models also demands billions of floating-point operations. Caching is a natural solution to reduce LLM inference costs on repeated queries. However, existing caching methods are incapable of finding semantic similarities among LLM queries, leading to unacceptable false hit-and-miss rates.
This paper introduces MeanCache, a semantic cache for LLMs that identifies semantically similar queries to determine cache hit or miss. Using MeanCache, the response to a user’s semantically similar query can be retrieved from a local cache rather than re-querying the LLM, thus reducing costs, service provider load, and environmental impact. Existing LLM Cache trains a single model at the server to identify similarities across multiple queries, thus suffering from privacy, cache scalability, and wasteful query requests. MeanCache leverages Federated Learning (FL) to collaboratively train a query similarity model in a distributed manner across numerous users without violating privacy. By placing a local cache in each user’s device and using FL, MeanCache reduces the latency and costs and enhances model performance, resulting in lower cache false hit rates. MeanCache compresses the embedding dimensions to minimize cache storage and also finds the optimal cosine similarity threshold. Our experiments, benchmarked against the GPTCache, reveal that MeanCache attains an approximately 17% higher F-score and a 20% increase in precision during semantic cache hit-and-miss decisions. Furthermore, MeanCache reduces the storage requirement by 83% and accelerates semantic cache hit-and-miss decisions by 11%, while still surpassing GPTCache.
Federated Learning, Cache, Hit, Miss, Rate, ChatGPT, Llama, Semantic, Search, Large Language Models, LLMs, GPT, Machine Learning, Privacy, Embeddings, Cosine Similarity, Client
1. Introduction
Large Language Models (LLMs) like ChatGPT (cha, [n.d.]), Google Bard (bar, [n.d.]), Claude (cla, [n.d.]), and Llama 2 (Touvron et al., 2023) have demonstrated remarkable capabilities in understanding and generating human language, leading to significant advancements in a variety of applications ranging from search engines to conversational agents. The integration of LLMs into platforms like the Perplexity AI search engine  (per, [n.d.]b), Rabbit OS (rab, [n.d.]), Arc browser (arc, [n.d.]), and Bing Chat (bin, [n.d.]) has underscored their growing influence. For instance, in 2023, 1.9 billion queries were sent to Bing chat, reflecting the massive use of such systems (bin, [n.d.]).
Motivation. Traditional search engines respond to user queries with ranked result pages based on keyword matching (Lempel and Moran, 2003; Jansen et al., 2000; Markatos, 2001). In contrast, LLM-powered services like Perplexity AI search engine, Bing chat, and ChatGPT provide a single precise answer to a user query. However, generating these responses with LLMs, such as GPT-3, requires substantial computations and poses environmental challenges (Frantar et al., 2023; Wang et al., 2023; Tseng et al., 2024; Sachdeva et al., 2024; Ma et al., 2024; Strubell et al., 2019). For example, GPT-3’s 175B parameters in float16 format consume 326 GB memory, exceeding single GPU capacities and necessitating multi-GPU setups (Frantar et al., 2023). These requirements lead to high operational costs. Google Bard’s costs, for instance, can be ten times those of standard keyword searches, potentially adding an extra \$6 billion for half of Google’s search queries (AIi, [n.d.]). Consequently, LLM-based services charge users and limit query rates (per, [n.d.]a; ope, [n.d.]).
Users frequently submit repeated queries to web services (Lempel and Moran, 2003; Xie and O’Hallaron, 2002), with approximately 33% of search engine queries being resubmitted (Markatos, 2001). Caching serves as an effective technique in web services to address these duplicate queries, avoiding redundant computations, significantly improving response time, reducing the load on query processors, and enhancing network bandwidth utilization (Markatos, 2001; Saraiva et al., 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003; Brin and Page, 1998; Podlipnig and Böszörmenyi, 2003; Baeza-Yates et al., 2007). Applying caching to serve repeated queries in LLMs based web services can have a substantial impact, reducing billions of floating point operations, thereby decreasing operational costs and environmental impacts.
Problem. Existing keyword matching caching techniques  (Saraiva et al., 2001; Markatos, 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003) will fail to capture the semantic similarity among repeated queries to LLM based web services and will result in a significantly low hit rate. For instance, if a user sends the query ”How can I increase the battery life of my smartphone?” and later sends a similar query, ”Tips for extending the duration of my phone’s power source.”, the existing caches will not detect the semantic similarity between these two queries and will lead to a cache miss.
Zhu et al. (2023) and GPTCache(Bang, 2023) present server-side semantic caching for the LLMs based services to address the limitations of keyword-matching caching techniques. They determine if a new query is semantically similar to any query in the cache by using an embedding model. For instance, as per study in (Zhu et al., 2023), if a similar query is found, the server returns the response from the cache. Otherwise, a model multiplexer is used to select the most suitable LLM for the query to generate the response.
These semantic caches have several limitations. (1) They demand a significantly large central cache to store the queries and responses of all users, which violates users privacy. (2) They still incur the network cost of sending a user query to the server even if there is a cache hit. (3) The server-side embedding model to find the semantic similarity among queries is generic and is not tailored to each user’s querying patterns (i.e., model does not adapt to the behavior of each user). (4) Lastly, an end user will still be charged for the query even if the query is served from the server cache without sending it to the LLM.
Key Insights and Contributions of MeanCache. Our work, MeanCache, introduces a novel user-centric semantic caching system that tackles various limitations and challenges inherent in semantic search caches today. An effective semantic search cache must be efficient, compact, and privacy-preserving. MeanCache provides a privacy-preserving caching system that can return the response to repeated queries directly from the local cache of the user, bypassing the need to query the LLM based web service. MeanCache’s achieves these goals in the following ways:
The embedding model used to compute a query’s embedding vector should be compact and efficient, enabling swift computation of embeddings. MeanCache achieves this by employing a smaller transformer model, MPNet (Song et al., 2020), to efficiently compute the user’s query embeddings. MPNet’s smaller size and faster operation, compared to larger models like Llama 2 (Touvron et al., 2023), make it compatible with resource-limited user devices, thereby enabling a resource-efficient semantic cache. Our insight is that a smaller model can achieve performance comparable to larger models on custom tasks  (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). To customize the embedding model, MeanCache uses federated learning for training on user data without storing the data on the service provider. This privacy-preserving training not only customizes the embedding model to the user’s querying patterns but also enhances the performance (i.e., accuracy) of semantic caching without compromising user privacy (i.e., without storing user-data at the web server). The runtime performance of MeanCache is primarily influenced by the time taken to match a new query embedding vector with existing ones in the cache to find a semantically similar query. The search time is directly proportional to the dimensions of the embedding vector. To optimize runtime performance, MeanCache compresses the embedding vector by leveraging principal component analysis (PCA) (Pearson, 1901; Hotelling, 1933; Jolliffe and Cadima, 2016), effectively reducing the size of the embedding vector (i.e., projecting it to lower dimensional space).
In MeanCache, embeddings are computed for new queries and their similarity to cached queries is measured using cosine similarity, an effective metric for capturing semantic similarity between texts. However, determining an optimal threshold for cosine similarity is crucial for accurately identifying true cache hits and misses. For example, a low threshold value may result in a high false hit rate, incorrectly matching a new query with cached ones. MeanCache mitigates this by identifying the optimal threshold from the user’s cached queries, specifically the one that maximizes the F score, thereby minimizing the false hit-and-miss rates on the client’s local data. The cost of finding this optimal threshold is approximately equivalent to one epoch of training on the client’s local data. It’s important to note that the optimal threshold, along with the local weights of the client’s embedding model (post local training), is sent to the central server to update the federated learning global model and the global cosine similarity threshold. This global threshold proves beneficial for similarity matching for new users who do not have cached queries. When a new query matches a cached one on a given cosine similarity threshold, MeanCache results in cache hits and delivers responses from the local cache. If not, it processes cache misses by querying the service provider and storing new query-response pairs for future use.
Evaluations. We compare MeanCache with GPTCache(Bang, 2023), a widely-used open-source semantic cache for LLM-based web services. GPTCache (Bang, 2023) is the most closely related work to MeanCache and has garnered over 6,000 stars on GitHub(zil, [n.d.]). We benchmark MeanCache’s performance against GPTCache on the GPTCache’s dataset (dat, [n.d.]) to demonstrate its effectiveness and highlight the improvements MeanCache offers over existing solutions. Our experiments utilize the Flower FL framework (Beutel et al., 2020) and the SBERT library(Reimers and Gurevych, 2019) to train embedding models in a distributed and privacy-preserving manner with 20 clients. The client size is constrained by the limited size of the GPTCache dataset, which is insufficient to distribute among hundreds of clients. However, our findings are not dependent on the number of clients, and MeanCache’s evaluation setup aligns with the prevailing evaluation standard for new studies in FL (Avdiukhin and Kasiviswanathan, 2021; Wang et al., 2020a). MeanCache surpasses GPTCache (Bang, 2023) by achieving a 17% higher F score and approximately a 20% increase in precision in end-to-end deployment for identifying duplicate queries to LLM-based web services. MeanCache’s embedding compression utility approximately reduces storage and memory needs by 83% and results in 11% faster semantic matching, while still outperforming the state-of-the-art GPTCache.
In summary, MeanCache introduces the design of the first user-centric semantic cache that utilizes federated learning. This cache serves repeated queries to LLM-based web services directly from a user’s local cache, thereby eliminating the need to query web services. Users’ queries are not shared with service providers or stored on centralized servers. MeanCache effectively enhances the embedding model to generate high-quality query embeddings using federated learning. This improvement not only boosts MeanCache’s accuracy, exemplified by an increased true hit rate, but also conserves the user’s query quota and reduces the redundant computations required by LLM services, all without compromising user privacy.
2. Background
2.1. Federated Learning
Federated Learning is a distributed training of a machine learning (ML) model. The key benefit of this approach is that users’ data never leaves its source. It mainly consists of four steps. In the first step, the central server shares the same model with all the participating clients. In the second step, each client trains the model on its local data. In the third step, the clients send the updated model to the central server. In the fourth step, the central server aggregates the received models’ weights to form a new global model. The aggregation is done mainly by averaging the weights of the models, and there are different aggregating/fusion algorithms, such as FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), and FedMA (Wang et al., 2020b). The FedAvg algorithm is the most popular one, and it aggregates the models with the following equation:
(1) $W_{global}^{t+1}=\sum_{k=1}^{K}\frac{n_{k}}{n}w_{k,t}$
In the above equation, $W_{global}^{t+1}$represents the new global model, $w_{k,t}$ denotes the model of the $k^{th}$ client at the $t^{th}$ round, $n_{k}$ are the number of samples of the $k^{th}$ client, and $n$ is the total number of samples of clients participating in the given round. Steps 1-4 constitute a single Federated Learning training round. This procedure is reiterated for multiple rounds until the model reaches convergence.
2.2. Transformer Architecture
The transformer architecture is a deep learning model based on the attention mechanism, which allows the model to focus on the important parts of the input (Vaswani et al., 2017). The transformer architecture is primarily used for sequence-to-sequence tasks in NLP, such as machine translation, text summarization, and question-answering. It consists of two main components: the encoder and the decoder. The encoder takes the input sequence and encodes it into a vector representation. The decoder takes the encoded vector representation and decodes it into the output sequence. Both the encoder and the decoder are composed of multiple layers. Each layer in the transformer architecture has two sub-layers: the multi-head attention layer and the feed-forward layer. The multi-head attention layer performs the attention mechanism, which allows the model to focus on different parts of the input sequence. The attention mechanism calculates the importance of each input element by considering its relationship with other elements in the sequence. This enables the model to capture long-range dependencies and improve performance on tasks that require an understanding of context. The feed-forward layer is a fully connected layer that applies a non-linear transformation to the output of the multi-head attention layer. It helps the model capture complex patterns and relationships in the data. The transformer architecture also incorporates positional encoding, which provides information about the position of each element in the input sequence. This helps the model understand the order of the elements, which is crucial for tasks like machine translation.
2.2.1. Embeddings and Cosine Similarity
Transformer model can encode the given text into embeddings. Embeddings represent words or phrases as dense vectors in a high-dimensional space and capture the semantic meaning of words or phrases, and therefore, words or sentences with similar meanings will have similar vector representations (Reimers and Gurevych, 2019; Karpukhin et al., 2020; Ni et al., 2022). For example, the words ”king” and ”queen” are semantically similar because they both refer to royalty. Therefore, their embeddings will be close to each other in the vector space. These embeddings encode the contextual information and semantic meaning of the sentences. Cosine similarity is a measure of similarity between two vectors. In the context of embeddings, cosine similarity is often used to quantify the similarity between sentence embeddings. It is calculated by taking the cosine of the angle between the two vectors. The cosine similarity ranges from -1 to 1, where -1 indicates completely opposite vectors, 0 indicates orthogonal (perpendicular) vectors, and 1 indicates identical vectors. Cosine similarity ($\theta$) between embeddings E1 and E2 is measured as:
(2) $\theta=\frac{{\textbf{E1}}\cdot{\textbf{E2}}}{\|{\textbf{E1}}\|\|{\textbf{E2}}\|}$
Here, E1 and E2 represent the two word embeddings, ”dot” represents the dot product of the vectors, and ”——E1——” and ”——E2——” represent the Euclidean norms (magnitudes) of the vectors.
3. MeanCache-A User-Centric Semantic Caching System for LLMS
Caching is a well-known technique for improving the performance of content serving systems (Markatos, 2001; Saraiva et al., 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003). Caching at the end-user level can save computation time, and similar repeated queries can be served from the local cache. However, designing a user-centric semantic caching system for the users who engage with LLM-assisted web services, such as Perplexity AI, Bing and ChatGPT, is not trivial and poses several challenges and many factors to consider. Following are the challenges and factors to consider while designing a caching system for LLMs.
3.1. Research Challenges
Web server caching techniques (Baeza-Yates and Saint-Jean, 2003; Long and Suel, 2005; Saraiva et al., 2001; Lempel and Moran, 2003; Marin et al., 2010) primarily utilize keyword matching, returning a list of links that align with the user’s query. However, these techniques fall short when compared to Language Learning Model (LLM) based systems like ChatGPT or the Perplexity AI search engine, which provide a unified response. This discrepancy becomes evident when a user sends similar but not identical queries. For example, queries like ”Draw a line-plot in Python?” and ”Plot a linear graph using Python” may yield suboptimal results, leading to false misses.
GPTCache(Bang, 2023) and Zhu et al. (2023) propose server-side semantic caching for LLM-based services to return a response from the cache if the query is semantically similar to any query in the cache. This approach reduces the computational cost of LLMs. However, these techniques primarily designed for server-side and store users’ queries on the server which is privacy risk. Furthermore, a user will still be charged for the query even if the query is served from the cache. Additionally, GPTCache(Bang, 2023) also propose Llama 2 to generate superior embeddings, thereby improving semantic matching accuracy. However, this proposal has several limitations. Large language models like Llama 2 are not only sizeable, being gigabytes in size, but they also demand substantial computational resources to generate embeddings. The deployment of such models is impractical due to their size and significant computational overhead for semantic matching. Moreover, the large dimension of Llama 2’s embeddings (Llama 2 embeddings dimension is 4096) can cause considerable overhead when matching a new user query embeddings with cached queries embeddings, as the search time is proportional to the dimensions of the embedding vector. Additionally, high-dimensional embeddings also utilize more memory and storage (e.g., the embeddings generated by Llama 2 for a single query requires approximately 32.05 KBs). An alternative could be to use a smaller pre-trained embedding model on the user side. However, these models may not offer the same level of accuracy and are unable to utilize the vast amount of distributed data available on the user’s device. Storing user queries at the central server and fine-tuning smaller embedding model raises privacy concerns as the user queries are stored at the central server.
Even having optimal embedding model, the task of determining whether a new query embeddings are similar to locally cached queries embeddings is challenging. After generating query embeddings using a transformer model, a similarity metric such as cosine similarity is used to ascertain if the new query embeddings match the cached embeddings of previous queries. This process involves setting a threshold for cosine similarity, which is a delicate balance. A low threshold may result in false positives, matching queries to irrelevant cached queries, while a high threshold could lead to false negatives, missing too many real matches. Therefore, setting an appropriate threshold is both mandatory and challenging.
In conclusion, an efficient user-centric semantic cache system should operate on the user side, utilizing a transformer embeddings model with low inference time to generate embeddings. It should autonomously determine the optimal threshold for cosine similarity to enhance the accuracy of embeddings matching to find similar queries. Additionally, the semantic search time of finding similar queries from the cache should be optimized. Ultimately, this caching system should leverage the vast amount of distributed data available on the user’s device to improve embeddings generation, while preserving user privacy. This approach will enhance the accuracy of semantic matching, reducing both false misses and false hits.
3.2. MeanCache’s System Design
MeanCache is a user-centric semantic caching system, optimized to function on the user side. It enhances the performance of the semantic caching system by leveraging the collective intelligence of multiple users. The system design of MeanCache is briefly described as follows: (1) MeanCache employs a small embedding model, which has a lower computational overhead compared to large Language Learning Model (LLM) based embedding models, such as Llama 2. (2) MeanCache uses federated learning for training to fine-tune the embedding model. This approach generates high-quality embeddings and improves the accuracy of embedding matching for retrieving similar queries. Notably, federated learning does not necessitate the storage of user queries in a central location, as it can train a high-quality model using the users’ distributed data. (3) To reduce storage and memory overhead, as well as to expedite the search time for finding similar queries in the cache, MeanCache compresses the embeddings using Principal Component Analysis (PCA). (4) MeanCache autonomously determines the optimal threshold for cosine similarity. It does this by varying the threshold and selecting the one that optimizes the F-score of the cache. Next we describe the system design of MeanCache in detail.
3.2.1. Privacy-Preserving Embedding Model FL Training
FL training is a decentralized approach to train machine learning models. It allows the users/clients to train the embedding model locally and share the updated weights with the server. The server aggregates the updated weights from multiple users, evaluates the global model, and redistributes the updated model to the users. This approach ensures that the user’s privacy is maintained, and the collective intelligence of multiple users is leveraged to improve the performance of the caching system. In the first step, the server sends the initial weights of the embedding model ($W_{global}^{t+1}$) to a subset of users. The subset of users are randomly selected or selected based on their battery level, network bandwidth, or performance history. Additionally, the server also sends the hyperparameters (e.g., learning rate, batch size, epochs) necessary for FL training of the embedding model.
Client Training.
Upon receiving the embedding model ($W_{global}^{t+1}$) from the server, each client replaces its local embedding model weights with the newly received weights. Subsequently, each client proceeds to train the embedding model locally on its unique local dataset. Local training employs a multitask learning approach, integrating two distinct loss functions: contrastive loss  (Reimers and Gurevych, 2019) and multiple-negatives ranking loss  (Henderson et al., 2017; Reimers and Gurevych, 2019). These loss functions update the weights of the embedding model during the training process. The contrastive loss function operates by distancing unique (non-duplicate) queries from each other within the embedding space, thereby facilitating the differentiation between duplicate and non-duplicate queries. Unlike contrastive loss, the multiple-negatives ranking loss function minimizes the distance between positive pairs (duplicate queries) amidst a large set of potential candidates i.e., multiple-negatives ranking loss does not concentrate on distancing unique queries and its primary objective is to draw positive pairs (similar queries) closer within the embedding space.
This multitasks learning approach enables MeanCache to adjust to diverse query patterns exhibited by users. For instance, some users may generate more repetitive queries compared to others, while certain users may not produce any repetitive queries at all. Interestingly, MeanCache’s multitask learning objective can benefit from learning even from a user with no repetitive queries. This is because MeanCache’s global embedding model ($W_{global}^{t+1}$) will learn to widen the distance between unique queries, thereby effectively learning true misses of the non-duplicate queries and consequently minimizing the false hits during the search process.
Finding the Optimal Threshold for Cosine Similarity.
In addition to privacy preserving training of the embedding model, identifying the optimal threshold ($\tau$) for cosine similarity is equally important. The range of $\tau$ is between 0 and 1. This threshold ($\tau$) dictates the level of similarity above which a cached query is considered relevant to the current user query. Setting the threshold too low could result in numerous false hits, leading to the retrieval of irrelevant queries from the cache. Conversely, a threshold set too high might cause many false miss, where relevant queries are not retrieved from the cache. Hence, finding the optimal threshold ($\tau$) for cosine similarity is vital.
After the client’s local training, MeanCache automatically determines this optimal threshold ($\tau$) from the clients’ data. MeanCache’s varies the threshold $\tau$ to find the optimal threshold that optimizes the F-score of the cache (Section 4.2). The client/user will use the optimal threshold to find similar queries in the cache. By finding the optimal threshold, MeanCache effectively balances between true hits and true misses, therefore yielding improved accuracy in semantic similarity matching to return the response from the cache on duplicate queries.
Aggregation.
After client local training and finding the optimal threshold, each client sends updated weights of the global model ($W_{global}^{t+1}$) and optimal threshold ($\tau$) to the server. The server aggregates the updated weights from multiple users to form a new embedding model ($W_{global}^{t+1}$) using FedAvg (McMahan et al., 2017). The server also computes the mean of the received optimal thresholds from the clients for a global optimal threshold ($\tau_{global}$). The benefit of finding $\tau_{global}$ is that when a new user joins the system, the user will not have queries to find its own optimal threshold. In such cases, the system can use $\tau_{global}$ as a starting point for semantic matching.
After finding the global optimal threshold and the global embedding model, the server then redistributes the updated embedding model to the users for the next round of FL training. This process is repeated until the semantic matching accuracy of the MeanCache meets a certain performance threshold (i.e., false hits and false misses are minimized). This approach ensures that the user’s privacy is maintained (i.e., clients queries are never stored at the central server), and the collective intelligence of multiple users is leveraged to improve the performance of the MeanCache. After the completion of the FL training, each client will have access to a fin-tuned embedding model ($W_{global}^{t+1}$) to generate high-quality embeddings which can capture the complex semantics of a user query. For instance, if previously the embedding model was unable to differentiate between ”I love cats” and I do not love cats”, the updated embedding model will be able to differentiate between such queries as it may learn such different semantics from the other users.
3.2.2. Embeddings Compression uisg PCA
MeanCache’s search time performance directly depends on the dimensionality of the embeddings generated by the embedding model. High-dimensional embeddings not only require more memory and storage but also increase the search time for finding similar queries in the cache. This is because the search time is proportional to the dimensions of the embedding vector. Computing the cosine similarity between two high-dimensional embeddings (i.e., new query embeddings and with each embedding in the cache) is computationally expensive and time-consuming. Thus, to improve the search time for finding similar queries in the cache, it is essential to reduce the dimensionality of the embeddings with minimal impact on MeanCache’s performance.
Principal Component Analysis (PCA) is a dimensionality reduction technique that is widely used to compress high-dimensional data into a lower-dimensional space  (Pearson, 1901; Hotelling, 1933; Jolliffe and Cadima, 2016), while still maintaining the most important information. First, MeanCache generates embeddings for all the users’ queries using the embedding model. Next, MeanCache applies PCA to learn the principal components of all the queries embeddings generated in the previous step. MeanCache integrates the learned principal components as an additional layer in the embedding model. This new layer will project the original embeddings onto the lower dimensional space, producing compressed embeddings.
Now, when a non-duplicate query is received, MeanCache will use the updated embedding model (with PCA layer) to generate the compressed embeddings for the new query and store the query, response and the compressed embedding in the cache. Storing the compressed embeddings in the cache will significantly reduce the storage and memory overhead of the embeddings. Next, when a duplicate query is received, MeanCache will use the same embedding model with PCA components to generate the compressed embeddings for this duplicate query and find the similar queries in the cache. Since the embeddings are compressed, the search time for finding similar queries in the cache will be significantly reduced.
4. Evaluation
We conduct evaluations of MeanCache against the baseline GPTCache (Bang, 2023) to demonstrate that MeanCache achieves optimal performance while preserving user-privacy (i.e., without storing the user queries at the server). For a fair comparison between MeanCache and GPTCache, we employ the optimal configuration as described in the GPTCache study (Bang, 2023). This configuration utilizes Albert (Lan et al., 2019) and sets the cosine similarity threshold at 0.7 to determine the cache hit or miss.
4.0.1. Transformer Models and Datasets.
For extensive evaluations of MeanCache, we utilize the Llama 2 (Touvron et al., 2023), MPNet (Song et al., 2020), and Albert (Lan et al., 2019) transformer models to generate embeddings.
We evaluate MeanCache using the GPTCache dataset. The dataset is partitioned into training, testing, and validation subsets. The training and validation datasets are randomly distributed among the clients, with each client receiving non-overlapping data points. During local training, each client utilizes its training dataset to update its local embedding model and employs the validation dataset to determine the optimal threshold for cosine similarity (Section 4.2). The testing dataset, located at the central server, facilitates a fair comparison between MeanCache and GPTCache (Bang, 2023). It is important to note that usually in real-world deployments, the test dataset may not be available at the central server, and the metrics reported by clients are used to evaluate the performance of the global model.
4.0.2. Experimental Setup
The experiments are conducted on a high-performance computing cluster, equipped with 128 cores, 504 GB of memory, and four A100 Nvidia GPUs, each with 80 GB of memory. We utilize the Flower FL (Beutel et al., 2020) library to simulate a federated learning setup. Additionally, the SBERT (Reimers and Gurevych, 2019) library is employed to train the embedding model on each client. The number of clients participating in FL training are 20. The number of clients is restricted due to the limited size of the GPTCache dataset, which is inadequate for distribution among hundreds of clients. However, we believe MeanCache results are not influenced by the number of clients, and the evaluation setup of MeanCache is consistent with the evaluation standard in FL (Avdiukhin and Kasiviswanathan, 2021; Wang et al., 2020a).
4.0.3. Evaluation Metrics
In caching systems, the efficacy has traditionally been gauged by cache hit-and-miss rates. A cache hit implies the data or query is retrieved from the cache, whereas a cache miss indicates the opposite. Semantic caching introduces a nuanced classification: true and false hits, alongside true and false misses. A true hit signifies a correct match between a query and a similar cached query, whereas a false hit is an incorrect match with a non-similar cached query.
Thus, semantic caching complexity renders traditional hit/miss metrics potentially misleading. For example, a query might incorrectly match with an irrelevant cached query (deemed a hit traditionally) due to semantic matching. This necessitates more sophisticated metrics to accurately assess a semantic cache’s performance. We adopt precision, recall, F1 score, and accuracy, for a comprehensive evaluation of MeanCache against GPTCache. These metrics are defined as follows:
Precision. The ratio of true positive hits to all positive hits (including both true positives and false positives). In semantic caching, this measures how many of the queries matched to a cached query are correctly matched. $\text{Precision}=\frac{TP}{TP+FP}$ where $TP$ represents true positives (true hits) and $FP$ represents false positives (false hits).
Recall. The ratio of true positive hits to all relevant items (including both true positives and false negatives). In semantic caching, this assesses the proportion of correctly matched queries out of all queries that should have been matched to a cached query. $\text{Recall}=\frac{TP}{TP+FN}$ where $FN$ represents false negatives (false misses).
[…]
$F_{\beta}=(1+\beta^{2})\cdot\frac{\text{Precision}\times\text{Recall}}{(\beta^{2}\times\text{Precision})+\text{Recall}}$
Accuracy. The ratio of correctly identified queries (both true hits and true misses) to all queries. $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ where $TN$ represents true negatives (true misses).
4.0.4. Research Questions
  • Is it possible to train an embedding model in a privacy-preserving manner without the need for centralized data?
  • Besides preserving user query privacy, how does MeanCache perform in comparison to GPTCache in terms of performance metrics?
  • What effect does the cosine similarity threshold have on the performance of MeanCache?
  • What level of overhead does MeanCache impose on the user side? Is it possible to reduce the embedding dimension to save storage space while still outperforming the baseline?
  • GPTCache (Bang, 2023) suggests using Llama 2 to generate embeddings to improve semantic matching. Is it feasible to use Llama 2 to compute embeddings at the user side for semantic matching?
01020304050FL Training Rounds0.750.800.850.900.95ScoreF1PrecisionRecallAccuracy
Figure 1. FL training enhances MPNet’s ability to generate high-quality embeddings by leveraging users’ private data, thereby improving MeanCache’s query matching precision from 0.74 to 0.85, demonstrating its effectiveness.
01020304050FL Training Rounds0.750.800.850.900.95ScoreF1PrecisionRecallAccuracy
Figure 2. FL training boosts MeanCache’s user query matching precision with Albert embedding model from 0.74 to 0.81, ensuring privacy.
4.1. Privacy Preserving Embeddings Model Training
Storing user or client queries on the server side presents a potential privacy risk. To address this, each client can retain its local data on its own device. The ensuing challenge is how to train an embedding model that can also utilize the distributed data from all clients. Federated Learning is recognized for training neural networks in a privacy-preserving manner. As such, MeanCache employs FL to train and fine-tune an embedding model, thereby preserving privacy and leveraging the dataset residing on the client’s side. In this section, our objective is to evaluate whether FL training can progressively enhance the embedding model to generate high-quality embeddings for user queries. To simulate this scenario, we distribute the training dataset among 20 clients. In each round, we sample 4 clients, conducting a total of 50 FL training rounds. Each client trains its embedding model for 6 epochs, operating on a dedicated A100 GPU. We conduct two experiments with the Albert and MPNet models. The batch size is set to 256 for the Albert model, and for MPNet, it is set to 128 during the local training by participating clients.
Figures 1 and 2 depict the performance of MeanCache as FL training progresses. The X-axis represents the training round, while the Y-axis shows the performance metrics such as F-score, precision, recall, and accuracy of the global model ($W_{global}^{t+1}$). As illustrated in Figure 1, the F-score for MPNet increases from 0.82 to 0.88, and for Albert, it rises from 0.83 to 0.86, as shown in Figure 2. Similarly, precision for MPNet significantly increases from 0.74 to 0.85, as depicted in Figure 1, and for Albert, it increases from 0.74 to 0.81, as demonstrated in Figure 2. Given that MPNet is a more robust transformer architecture compared to Albert, it is also observed during our training that MPNet outperforms Albert, exhibiting superior learning in FL settings.
Summary. Overall, FL training increases 11% precision of MeanCache for MPNet and a 7% increase for Albert. Therefore, it is evident that the performance of the embedding model to generate high-quality embeddings can improve in a privacy-preserving manner using FL training, that is, without storing user queries at the central server.
0.00.20.40.60.81.0Threshold0.250.500.751.00ScoreF1PrecisionRecallAccuracy
Figure 3. Extremes in threshold harm semantic matching; MeanCache autonomously optimizes MPNet’s threshold at 0.83 for an F score of 0.89.
0.00.20.40.60.81.0Threshold0.000.250.500.751.00ScoreF1PrecisionRecallAccuracy
Figure 4. For Albert embedding model, MeanCache identifies an optimal threshold of 0.78, achieving a peak F-score of 0.88 on validation data.
4.2. Cosine Similarity Threshold Impact on Semantic Matching
The process of semantic matching for a new user query begins by generating the embeddings of the user’s query ($E_{q}$) using the embedding model. Following this, the cosine similarity ($\theta$) is computed with the cached embeddings. If the cosine similarity $\theta$ exceeds the threshold $\tau$, the cache is hit and the response to the user query is returned from the cache. Therefore, the cosine similarity threshold $\tau$ plays a crucial role in determining the similarity between a user query and cached entries. A low threshold value of $\tau$ can lead to false hits (incorrect matches), while a high threshold might overlook the appropriate matches (i.e., false cache misses or false negatives). Hence, identifying an optimal threshold is essential.
To illustrate this, MeanCache varies the threshold $\tau$ from 0 to 9 and evaluates the performance metrics F-Score, precision, recall, and accuracy with an equal distribution of duplicate and non-duplicate queries from the validation data to avoid bias.
Figures 3 and 4 show how the cosine similarity threshold ($\tau$) affects MeanCache’s performance. The X-axis represents the threshold $\tau$ values, and the Y-axis denotes the performance metrics. For instance, at a 0.3 threshold, MeanCache’s semantic matching accuracy using MPNet is 57%, with a precision of 54% as shown in Figure 3. Similarly, with Albert at the same threshold, the accuracy is 55%, and precision is 53% (Figure 4). Precision typically improves with an increase in threshold. However, beyond a certain point, increasing the threshold $\tau$ leads to a decline in F score, accuracy, and recall, i.e., higher value of $\tau$ increases false cache misses (false negatives). Therefore, setting an optimal threshold is imperative.
For MPNet, the optimal threshold $\tau$ is identified at 0.83, achieving an F1 score of 0.89, precision of 0.92, and accuracy of 0.90 (Figure 3). For Albert, the optimal threshold is 0.78, with an F1 score of 0.88, precision of 0.86, and accuracy of 0.88. Note that in the rest of the evaluations, we use these thresholds for MPNet and Albert with MeanCache.
Summary. Experimental evaluations indicate that GPTCache’s suggested threshold of 0.7 will result in suboptimal performance during semantic matching. Moreover, the optimal threshold $\tau$ values varies with the embedding model. MeanCache optimally adjusts the threshold based on user data, outperforming GPTCache’s suggested threshold by 16% in precision and 4% in F score for MPNet, and by 10% in precision and 2% in F score for Albert. This highlights the importance of fine-tuning the cosine similarity threshold to its optimal value.
020406080100Query ID0.00.51.0Time (s)Llama 2Llama 2+GPTCacheLlama 2+MeanCache
Figure 5. Response times of 100 randomly sampled user queries to the Llama 2-based LLM service in three scenarios: without any semantic cache, with GPTCache, and with MeanCache. The integration of a semantic cache does not add significant overhead to non-duplicate queries, meaning it does not impede the performance of the LLM-based service. Moreover, it significantly reduces the average response times for duplicate queries (70-99) by serving them from the local cache.
020406080100Query IDHitMissLabelReal LabelGPTCache Predicted LabelMeanCache Predicted Label
Figure 6. Comparison of MeanCache and GPTCache on a set of 100 queries, including 70 non-duplicate and 30 duplicate queries, sent to the Llama 2-based LLM service. Queries 0 to 69 are non-duplicate (i.e., real label is miss), and GPTCache produces significantly higher false hits on these unique queries compared to MeanCache.
01Predicted Label01Real Label6118966234
(a) MeanCache
01Predicted Label01Real Label46723346254
(b) GPTCache
Figure 7. Confusion matrices for MeanCache and GPTCache for comprehensive insight. Total queries are 1000, comprising 700 unique and 300 duplicate queries. Among the 700 unique queries, MeanCache produces only 89 false hits, while GPTCache generates a significantly larger number of false hits, 233 in total. False hits are undesirable as they require the user to resend the query to the LLM service to obtain the correct response.
Metrics GPTCache MeanCache MeanCache
(MPNet) (Albert)
F score 0.56 0.73 0.68
Precision 0.52 0.72 0.66
Recall 0.85 0.78 0.77
Accuracy 0.72 0.85 0.81
Table 1. MeanCache outperforms GPTCache (Baseline)
100020003000(a) Number of Cached Queries02000400060008000Storage Size (KBs)100020003000(b) Number of Cached Queries0.010.020.03Avg. Search Time (s)100020003000(c) Number of Cached Queries0.50.60.7F ScoreGPTCacheMeanCache (MPNet)MeanCache (Albert)MeanCache-Compressed (MPNet)MeanCache-Compressed (Albert)
Figure 8. (a)MeanCache’s embedding compression utility significantly reduces storage space by 83% compared to GPTCache. (b) MeanCache’s semantic matching speed is 11% faster with compression enabled, while still outperforming GPTCache. (c) MeanCache’s F score is slightly lower with compression enabled, but it still outperforms GPTCache.
4.3. MeanCache Comparison with Baseline (GPTCache)
Markatos (2001); Xie and O’Hallaron (2002) have shown that users often send repeated queries to online services. While these repeated queries typically make up a smaller volume than the total query count, identifying them at the client-side can significantly reduce server load, decrease bandwidth requirements, and improve response times. Therefore, detecting repeated queries not only benefits LLM based services but also enhances the overall user experience and conserves the permissible query quota. We evaluate MeanCache against GPTCache, a baseline semantic cache, to assess improvements in user experience, precision, recall, and F score. We select a sample of 1000 queries, 30% of which are repeated queries (i.e., 300 queries are repeated), and load these 1000 queries as cached queries. Note that repeated queries are usually fewer than non-repeated queries. Thus, we use 30% as repeated queries, a similar percentage previously observed for web services (Markatos, 2001).
Initially, we send a new set of one thousand queries to Llama 2 (i.e., without any semantic cache) to establish a baseline for response times. We limit responses to 50 tokens to reflect practical response sizes, although actual sizes can be much larger. Note that MeanCache’s performance is not dependent on the response as it only matches the queries. Next, we send these queries to Llama 2 based local LLM service with MeanCache and GPTCache to measure the response times and performance metrics (e.g., precision, recall, F-score) respectively. An analysis of a random subset of 100 queries (70 non-duplicate and 30 duplicate) from the 1000 queries shows the impact of caching on response times in Figure 5 and cache hit and miss rates in Figure 6. The X-axis represents the query and the Y-axis represents each semantic cache hit and miss alongside the real label in Figure 6. The Y-axis in Figure 5 shows the response time of each query without any cache (Llama 2), with GPTCache (Llama 2+ GPTCache), and with MeanCache (Llama 2+MeanCache). Figure 5 demonstrates that implementing a semantic cache does not impede the performance (queries ranging from 0 to 69) and improves the user experience as response times reduces on duplicate queries (queries 70 to 99).
However, Figure 6 shows that GPTCache produces significant false hits on non-duplicate queries (queries 0 to 69) compared to MeanCache, detracting from user experience. Each false hit means the user receives an incorrect query response from the cache and must resend the same query to LLM service, resulting in a poor user experience. To prioritize precision, we adjust the beta value in the F score to 0.5, valuing precision twice as highly as recall to ensure user satisfaction by avoiding false positives. This decision is driven by the need to minimize user inconvenience caused by incorrect cache hits. A false miss (low recall) does not require user intervention as the false miss query will be automatically routed to the LLM. Thus, precision in semantic caching is more important than recall.
Table 1 and the confusion matrices in Figure 7 highlight MeanCache’s superior performance over GPTCache on the 1000 user queries. Notably, MeanCache with MPNet achieves a precision of 0.72, significantly surpassing GPTCache’s 0.52. This superiority is evident in the lower false positive rates (i.e., false hits) shown in Figure 6(a). The number of false hits for MeanCache is 89 (Figure 6(a)), while GPTCache has 233 false hits, as depicted in Figure 6(b). In practical terms, this means that with GPTCache, the end user has to manually resend 233 queries to the LLM service to get the correct responses, compared to only 89 queries with MeanCache. While GPTCache’s recall is higher than MeanCache’s, as we discussed earlier, precision is significantly more important than recall, and MeanCache outperforms GPTCache in this regard. Overall, the F-score of MeanCache with MPNet is 0.73 and 0.68 with the Albert embedding model, both of which outperform GPTCache’s F-score of 0.56.
Summary. In an end-to-end deployment with Llama 2 as LLM service, MeanCache significantly outperforms GPTCache. It demonstrates a 17% higher F score and a 20% increase in precision in optimal configuration. The substantial reduction in false cache hits enhances the end-user experience.
4.4. MeanCache Embedding Compression utility and Impact on Storage Space
Clients or users often face storage limitations compared to web servers. Storing embeddings in the local cache on the user side for semantic search demands memory storage. Various models yield embeddings with differing vector sizes; for example, the MPNet and Albert models produce an output embedding vector of 768 dimensions, whereas the Llama 2 model’s embeddings dimension size is 4096. The embedding vector size also affects semantic search speeds, where smaller vectors could enhance speed and lower resource demands.
Figure 8 illustrates the effects of MeanCache dimension reduction utility on the storage, semantic matching speed (overhead), and MeanCache’s performance (F score). The x-axis indicates the number of queries stored in the cache, while the y-axis shows storage size, average search time, and the F score in respective graphs. MeanCache-Compressed (MPNet) and MeanCache-Compressed (Albert) represent instances where MeanCache decreases the embedding dimensions from 768 to 64 by employing the compression, as detailed in MeanCache design (Section 3.2).
Figure 8 demonstrates that an increase in stored queries linearly raises storage needs. Yet MeanCache with compression enabled drastically lowers storage needs by 83% compared to GPTCache. Figure 8 also indicates that compression decreases the average search time, with MeanCache enabled compression being approximately 11% faster than without. Moreover, despite a slight decrease in F score with compression enabled, MeanCache still surpasses GPTCache. Furthermore, given the evidence from Figures 1 and 2 that MPNet produces more precise embeddings, and it is also clear from Figure 8 that MPNet’s embeddings are particularly resilient to compression and excel in semantic matching.
Summary. The application of embedding compression optimization in MeanCache offers substantial benefits, including an 83% savings in storage and an 11% faster search process, while still outperforming the established baseline (GPTCache).
Llama-2mpnetalbert0.000.020.04Computation Time (s)0.0400.0050.009Llama-2mpnetalbert0102030Storage Size (KBs)3266
Figure 9. Llama 2 takes significantly longer to compute embeddings and requires substantially more storage space than Albert and MPNet.
0.00.20.40.60.81.0Threshold0.000.250.500.751.00ScoreF1PrecisionRecallAccuracy
Figure 10. Llama 2 as an embedding model does not perform well in semantic matching, even at optimal thresholds. The maximum F1 score is 0.75, significantly lower than with the smaller embedding models (MPNet and Albert).
4.5. Infeasibility of Embedding Generation with Llama 2
GPTCache(Bang, 2023) recommend using Llama 2 for generating embeddings to enhance GPTCache’s performance. However, Llama 2’s embedding computation is expensive in terms of inference time, requires substantial storage, and incurs considerable overhead during semantic searches. For example, Llama 2 with its 7 billion parameters, demands 30 GB of memory (Siz, [n.d.]), whereas Albert and MPNet require only 43 MB and 420 MB, respectively.
To highlight the impracticality to generate query embeddings with Llama 2, we compare the embedding computation time, embedding storage space requirement of Llama 2, Albert and MPNet transformer models. Figure 9 shows the average time to compute the embeddings of a single query and storage requirements for the embeddings. Figure 9 shows that the average embedding computation time of Llama 2 is 0.04 seconds, while for Albert and MPNet the average computation time is 0.005 and 0.009 seconds respectively. Single query embeddings generated from Llama 2 takes approximately 32 KBs of space and embeddings generated by both MPNet and Albert only take 6 KBs, as shown in Figure 9.
Furthermore, we evaluate the performance of the Llama 2 on embedding generation and semantic matching. Figure 10 shows that the performance of Llama 2 with different cosine similarity threshold ($\tau$) and corresponding performance metric. We can see that the performance of Llama 2 is not good even with the optimal cosine similarity threshold, as also noted by researchers (Use, [n.d.]). The maximum F1 score achieved by Llama 2 is 0.75 which is quite low when compared with the optimal thresholds scores from the Figures 3 and Figure 4.
Smaller models tailored for specific tasks often surpass larger models in efficiency (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). Thus, diverging from GPTCache’s approach, we advocate for adopting smaller, more efficient embedding models for semantic caching. These models not only ensure optimal performance but also minimize the semantic cache’s overhead on users, featuring lower inference demands and reduced output embedding sizes, thereby facilitating deployment on edge devices.
Summary. Llama 2 as an embedding model is not a viable option for generating embeddings. Future enhancements might improve its performance to generate high-quality embeddings, yet the computational demands, semantic search duration, and storage requirements are likely to remain elevated, consequently impacting the system’s overall efficacy.
5. Related Work
Several caching systems are proposed to optimize the performance of search engines, databases and operating systems. Saraiva et al. (2001) suggest a two-tier dynamic caching architecture for web search engines to enhance response times in hierarchical systems. Utilizing LRU eviction at both levels, they demonstrate how the second-tier cache can significantly lower disk traffic and boost throughput.  Baeza-Yates and Saint-Jean (2003) propose a three-level index organization and  Long and Suel (2005) propose a three-tier caching. Xie and O’Hallaron (2002) examined two real search engine datasets to explore query locality, aiming to develop a caching strategy based on this concept. Their analysis centered on query frequency and distribution, assessing the feasibility of caching at various levels, such as server, proxy, and client side. Lempel and Moran (2003) proposed a novel caching technique, called Probability Driven Cache (PDC), to optimize the performance of search engines. PDC uses the probability of a query to be repeated to decide whether to cache the query or not. Fagni et al. (2006) proposes SDC (Static Dynamic Cache) that exploit the temporal and spatial locality present in the query stream to avoid processing the same query and thereby saving the computational resources. Baeza-Yates et al. (2007) explore efficient caching design for web search engines, comparing static and dynamic caching, and weighing caching query results against posting lists. Zhang et al. (2008) examines challenges in large search engines processing thousands of queries per second across vast document collections. The study focuses on index compression and caching optimizations, evaluating various inverted list compression algorithms, including novel variants, and analyzing different caching policies (e.g., LRU, LFU). Marin et al. (2010) introduces an innovative cache hierarchy aimed at boosting web search engines’ efficiency in processing user queries. By implementing a hierarchical cache system, the study aims to store different pieces of data useful for solving frequent queries, thereby enhancing overall query throughput and reducing single-user query latency and power consumption.
In the above studies, the caching systems are designed for traditional search engines that process keyword queries (i.e., key word matching) and return a list of links to the user as a response. However, when such caching systems are applied to LLMs based web servers or APIs, they will not be able to provide a single concise response and may return many false results. On top of that, such caching techniques will fail to capture the semantic similarity among repeated queries and will result in a significantly low hit rate. For instance, if a user sends the query ”How can I increase the battery life of my smartphone?” and later sends a similar query ”Tips for extending the duration of my phone’s power source.”, the existing caches will not be able to detect the semantic similarity between these two queries and will treat them as two different queries. Zhu et al. (2023) and (Bang, 2023) propose server-side caching for the services based on LLMs to reduce the massive computational cost of LLMs. For instance, Zhu et al. (2023) upon receiving a new query at the server checks if the query is semantically similar to any query in the cache. If the query is similar to any query in the cache then the server returns the response from the cache. Otherwise, a model multiplexer is used to select the most suitable LLM for the query.
Although, these techniques can handle the semantic similarity among the queries and can provide a single concise response, they raise privacy concerns as they involve storing user queries on the server side. Second, these techniques are static and unable to adapt to the behavior of each user. Furthermore, an end user will still be charged for the query even if the query is served from the cache. Thus, we need a user-centered semantic cache that can run on the user side that can benefit the user in terms of privacy, cost, and latency. Such a cache should be able to detect the semantic similarity among the queries and should be able to adapt to the behavior of each user in a privacy preserving manner. MeanCache provides aforementioned benefits without violating user privacy.
6. Conclusion
MeanCache is the first user-centric semantic cache for Language Model (LLM) based web services, such as ChatGPT. In MeanCache, clients train the global embedding model using federated learning on their local data, while preserving user privacy. The global model formed after the aggregation generates high-quality embeddings for semantic matching. When a new user’s query is duplicated, MeanCache semantically matches the new query with the user’s local cache and returns the most relevant results. This process reduces the computational cost of the LLM service, improves bandwidth and latency, and conserves the user’s query quota for the LLM service. Despite compressing embeddings and saving 83% of storage space, MeanCache outperforms the existing baseline semantic cache (GPTCache). Therefore, MeanCache presents a promising approach for training high-quality embedding models, preserving user privacy, and enhancing the performance of LLM-based web services by serving repeated queries from the user’s local cache with a significant low rate of false hits.